# Structure-Fair Quantum Circuit Complexity: An Auditable Information-Theoretic Lower Bound

### v2

@@ I.1 Core Setting
    Windowed Coarse-Graining: We adopt the conditional expectation maps $\{E_{\Xi}\}$ ($\Xi\geq 0$) from CWT [35], which correspond to a nested algebra of observables $\{A(\Xi)\}$ satisfying
    $E_{\Xi}(\sigma_{R})=\sigma_{R},\quad E_{\Xi}\circ E_{\Xi}=E_{\Xi},\quad\Xi_{1}\leq\Xi_{2}\Rightarrow A(\Xi_{1})\subseteq A(\Xi_{2})$ (I.1.1)
@@
    RCC Definitions: The information bandwidth $\Gamma_{R}$ is determined by $R$ and the hardware addressability (see Sec. 2.1). The standard and windowed RCC are:
    $C_{R}(\rho)=\frac{\log d_{R}-S(\rho)}{\log\Gamma_{R}},\quad C_{R,\Xi}(\rho)=\frac{\log d_{R}-S_{\Xi}(\rho)}{\log\Gamma_{R}}$ (I.1.2)
@@ I.2 Mathematical Properties of Windowed RCC
Proposition I.1 (Monotonicity and Limits). Under the setting (I.1.1), for any state $\rho$:
@@ I.3.1 Identity and Work Accounting in the Standard Case
$\Pi_{\Xi}=T_{\Xi}\cdot\log(\Gamma_{R})\cdot\left(\frac{\partial C_{R,\Xi}}{\partial\Xi}\right)_{E}$ (I.3.1)
Proof. From (I.1.2), we have $(\partial C_{R,\Xi}/\partial\Xi)_{E}=-(1/\log\Gamma_{R})\cdot(\partial S_{\Xi}/\partial\Xi)_{E}$. Using $\Pi_{\Xi}=-T_{\Xi}\cdot(\partial S_{\Xi}/\partial\Xi)_{E}$ then yields the result.
@@
$W_{\mathrm{info}}=\int_{\Xi_{i}}^{\Xi_{f}}\Pi_{\Xi}\mathrm{d}\Xi=\log\Gamma_{R}\cdot\int_{C_{R,\Xi_{i}}}^{C_{R,\Xi_{f}}}T_{\Xi}\mathrm{d}C_{R,\Xi}=\log\Gamma_{R}\cdot\langle T_{\Xi}\rangle_{C}\cdot\Delta C_{R,\Xi}$ (I.3.2)
@@ I.3.2 Generalization: Variable Bandwidth $\Gamma_{R}(\Xi)$
$\left(\frac{\partial C_{R,\Xi}}{\partial\Xi}\right)_{E}=-\frac{1}{\log\Gamma_{R}}\cdot\left(\frac{\partial S_{\Xi}}{\partial\Xi}\right)_{E}-\frac{C_{R,\Xi}}{\log\Gamma_{R}}\cdot\left(\frac{\partial\log\Gamma_{R}}{\partial\Xi}\right)$ (I.3.3)
@@
$\Pi_{\Xi}=T_{\Xi}\cdot\left[\log\Gamma_{R}\cdot\left(\frac{\partial C_{R,\Xi}}{\partial\Xi}\right)_{E}+C_{R,\Xi}\cdot\left(\frac{\partial\log\Gamma_{R}}{\partial\Xi}\right)\right]$ (I.3.4)
@@ I.4.1 Strict Lower Bound and Physical Form
$C^{*}(t_{f})-C^{*}(0)\leq\frac{2}{\hbar}\cdot t_{f}\cdot\langle\Pi_{\Xi}\rangle_{t}$ (I.4.1)
@@
$t_{f}\geq\frac{\hbar}{2\langle\Pi_{\Xi}\rangle_{t}}\cdot[C_{R}(\rho_{f})-C^{*}(0)]$ (I.4.2)
@@
$t_{f}\geq\frac{\hbar}{2\langle\Pi_{\Xi}\rangle_{t}}\cdot[C_{R}(\rho_{f})-C_{R}(\rho_{i})]$ (I.4.3)
@@
$t_{f}\gtrsim\frac{\hbar}{2\pi k_{B}\bar{T}_{\Xi}}\cdot\Delta C_{R}$ (I.4.4)
@@ I.4.2 Lower Bound with Full Correction Terms
$C_{R}^{\mathrm{full}}(\rho)=C_{R}(\rho)+\frac{\Delta_{\mathrm{spec}}(\rho)}{\log\Gamma_{R}}\\ C_{\log}(t,\varepsilon)=\frac{c_{1}}{\log\Gamma_{R}}\cdot\left[\log(\max\{1,C^{*}(t)\})+\log(1/\varepsilon)\right]$ (I.4.5)
@@
$t_{f}\geq\frac{\hbar}{2\langle\Pi_{\Xi}\rangle_{t}}\cdot[C_{R}^{\mathrm{full}}(\rho_{f})-C_{R}^{\mathrm{full}}(\rho_{i})-\Delta C_{\log}^{\mathrm{upper}}]$ (I.4.6)
@@ I.4.3 Boundary Case: $\Pi_{\Xi}$ Changing Sign
$t_{f}\geq\frac{\hbar}{2\langle\Pi_{\Xi}^{+}\rangle_{t}}\cdot[C_{R}(\rho_{f})-C^{*}(0)],\quad\langle\Pi_{\Xi}^{+}\rangle_{t}=\frac{1}{t_{f}}\int_{0}^{t_{f}}\Pi_{\Xi}^{+}(t)\mathrm{d}t$ (I.4.7)
@@ I.5.1 Handling of Support Set Leakage
If the physical state $\rho$ has a small leakage $p_{\mathrm{leak}}$ with respect to the target subspace $H_{R}$, the projected-and-renormalized state $\tilde{\rho}$ has a trace distance
@@
$|S_{\Xi}(\tilde{\rho})-S_{\Xi}(\rho)|\leq\delta\log(d_{R}-1)+h_{2}(\delta)$ (I.5.2)
@@
$|C_{R,\Xi}(\tilde{\rho})-C_{R,\Xi}(\rho)|\leq\frac{(\delta\log(d_{R}-1)+h_{2}(\delta))}{\log\Gamma_{R}}$ (I.5.3)
@@ I.5.2 Statistical Ledger for Operational Estimation
Following the protocol in Chapter 4, one can obtain a lower bound $L_{H}$ for $D_{H}^{\eta}(E_{\Xi}(\rho)||\sigma_{R})$ at a confidence level of $1-\delta$, which in turn gives a statistical lower bound for the main term of $C_{R,\Xi}$, $C_{R,\Xi}^{\mathrm{(main)}}\geq L_{H}/\log\Gamma_{R}$. Substituting this into (I.3.2) and (I.4.2) allows for the estimation of confidence intervals for $W_{\mathrm{info}}$ and $t_{f}$.

### v3

@@ I.1 Core Setting
    Windowed Coarse-Graining: We adopt the conditional expectation maps $\{E_{\Xi}\}$ ($\Xi\geq 0$) from CWT [liu2025heuristic], which correspond to a nested algebra of observables $\{A(\Xi)\}$ satisfying
    $E_{\Xi}(\sigma_{R})=\sigma_{R},\quad E_{\Xi}\circ E_{\Xi}=E_{\Xi},\quad\Xi_{1}\leq\Xi_{2}\Rightarrow A(\Xi_{1})\subseteq A(\Xi_{2})$ (97)
@@
    RCC Definitions: The information bandwidth $\Gamma_{R}$ is determined by $R$ and the hardware addressability (see II.1). The standard and windowed RCC are:
    $C_{R}(\rho)=\frac{\log d_{R}-S(\rho)}{\log\Gamma_{R}},\quad C_{R,\Xi}(\rho)=\frac{\log d_{R}-S_{\Xi}(\rho)}{\log\Gamma_{R}}$ (98)
@@ I.2 Mathematical Properties of Windowed RCC
Proposition I.1 (Monotonicity and Limits). Under the setting (97), for any state $\rho$:
@@ I.3.1 Identity and Work Accounting in the Standard Case
$\Pi_{\Xi}=T_{\Xi}\cdot\log(\Gamma_{R})\cdot\left(\frac{\partial C_{R,\Xi}}{\partial\Xi}\right)_{E}$ (99)
Proof. From (98), we have $(\partial C_{R,\Xi}/\partial\Xi)_{E}=-(1/\log\Gamma_{R})\cdot(\partial S_{\Xi}/\partial\Xi)_{E}$. Using $\Pi_{\Xi}=-T_{\Xi}\cdot(\partial S_{\Xi}/\partial\Xi)_{E}$ then yields the result.
@@
$W_{\mathrm{info}}$ $=\int_{\Xi_{i}}^{\Xi_{f}}\Pi_{\Xi}\,\mathrm{d}\Xi$ (100)
$=\log\Gamma_{R}\int_{C_{R,\Xi_{i}}}^{C_{R,\Xi_{f}}}T_{\Xi}\,\mathrm{d}C_{R,\Xi}$
$=\log\Gamma_{R}\,\langle T_{\Xi}\rangle_{C}\,\Delta C_{R,\Xi}$
@@ I.3.2 Generalization: Variable Bandwidth $\Gamma_{R}(\Xi)$
$\left(\frac{\partial C_{R,\Xi}}{\partial\Xi}\right)_{E}=-\frac{1}{\log\Gamma_{R}}\cdot\left(\frac{\partial S_{\Xi}}{\partial\Xi}\right)_{E}-\frac{C_{R,\Xi}}{\log\Gamma_{R}}\cdot\left(\frac{\partial\log\Gamma_{R}}{\partial\Xi}\right)$ (101)
@@
$\Pi_{\Xi}=T_{\Xi}\cdot\left[\log\Gamma_{R}\cdot\left(\frac{\partial C_{R,\Xi}}{\partial\Xi}\right)_{E}+C_{R,\Xi}\cdot\left(\frac{\partial\log\Gamma_{R}}{\partial\Xi}\right)\right]$ (102)
@@ I.4.1 Strict Lower Bound and Physical Form
$C^{*}(t_{f})-C^{*}(0)\leq\frac{2}{\hbar}\cdot t_{f}\cdot\langle\Pi_{\Xi}\rangle_{t}$ (103)
@@
$t_{f}\geq\frac{\hbar}{2\langle\Pi_{\Xi}\rangle_{t}}\cdot[C_{R}(\rho_{f})-C^{*}(0)]$ (104)
@@
$t_{f}\geq\frac{\hbar}{2\langle\Pi_{\Xi}\rangle_{t}}\cdot[C_{R}(\rho_{f})-C_{R}(\rho_{i})]$ (105)
@@
$t_{f}\gtrsim\frac{\hbar}{2\pi k_{B}\bar{T}_{\Xi}}\cdot\Delta C_{R}$ (106)
@@ I.4.2 Lower Bound with Full Correction Terms
$C_{R}^{\mathrm{full}}(\rho)=C_{R}(\rho)+\frac{\Delta_{\mathrm{spec}}(\rho)}{\log\Gamma_{R}}\\ C_{\log}(t,\varepsilon)=\frac{c_{1}}{\log\Gamma_{R}}\cdot\left[\log(\max\{1,C^{*}(t)\})+\log(1/\varepsilon)\right]$ (107)
@@
$t_{f}\geq\frac{\hbar}{2\langle\Pi_{\Xi}\rangle_{t}}\cdot[C_{R}^{\mathrm{full}}(\rho_{f})-C_{R}^{\mathrm{full}}(\rho_{i})-\Delta C_{\log}^{\mathrm{upper}}]$ (108)
@@ I.4.3 Boundary Case: $\Pi_{\Xi}$ Changing Sign
$t_{f}\geq\frac{\hbar}{2\langle\Pi_{\Xi}^{+}\rangle_{t}}\cdot[C_{R}(\rho_{f})-C^{*}(0)],\quad\langle\Pi_{\Xi}^{+}\rangle_{t}=\frac{1}{t_{f}}\int_{0}^{t_{f}}\Pi_{\Xi}^{+}(t)\mathrm{d}t$ (109)
@@ I.5.1 Handling of Support Set Leakage
If the physical state $\rho$ has a small leakage $p_{\mathrm{leak}}$ with respect to the target subspace $H_{R}$, the projected-and-renormalized state $\tilde{\rho}$ has a trace distance.
@@
$|S_{\Xi}(\tilde{\rho})-S_{\Xi}(\rho)|\leq\delta\log(d_{R}-1)+h_{2}(\delta)$ (111)
@@
$|C_{R,\Xi}(\tilde{\rho})-C_{R,\Xi}(\rho)|\leq\frac{(\delta\log(d_{R}-1)+h_{2}(\delta))}{\log\Gamma_{R}}$ (112)
@@ I.5.2 Statistical Ledger for Operational Estimation
Following the protocol Sec. IV, one can obtain a lower bound $L_{H}$ for $D_{H}^{\eta}(E_{\Xi}(\rho)||\sigma_{R})$ at a confidence level of $1-\delta$, which in turn gives a statistical lower bound for the main term of $C_{R,\Xi}$, $C_{R,\Xi}^{\mathrm{(main)}}\geq L_{H}/\log\Gamma_{R}$. Substituting this into (100) and (104) allows for the estimation of confidence intervals for $W_{\mathrm{info}}$ and $t_{f}$.